Actions
Commissioning and Ops Scheduling meeting : 26/1/2016¶
To be held in CR10 in the ATLAS Building @ 15:30 pm
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Agenda:¶
- Review of tasks and activities identified at last meeting
- Realignment work
- Action Item (Colin to prod John C) : need to get John's data with fiducials in it to get to Tony Middleton. Expected yesterday
- Simulation of material in the interspace
- Action Item : John to give these files to the tracker group to make certain tracker resolution doesn't change
- Action Item : John to do a simulation with air in the interspace and He in the absorber volume
- Action Item : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded
- Action Item : Colin will check with Alan Bross whether the LiH absorber can withstand vacuum
- Cooling water discussion
- Action Item : Colin and John to talk and agree on the plan. Discussion complete but further work needed
- ISIS Monitoring of SS magnets
- Action Item : Josef/Colin to check with Alan about monitoring requirements for the SS magnets.
- BLOCS and SOCS
- Action Item :
Steve to remind Adam that we need SOCs.
- Action Item :
- DS PS Installation
- Action Item : Colin is to get an estimate for power supply installation time. 2 sets of 2-3 days over a space of 2 weeks
- Realignment work
- Upcoming and ongoing activities : chaired by DC
- Operations calendar
- C&M Network reconfiguration
- Realignment work schedule
- Tracker patch panel leak work
- Date of DSA closure and work required before hand
- Henry's January workplan
- Daresbury workplan
- AOB
- Date of next meeting : February 2nd
Minutes¶
Present: Andy, Vicky Bayliss, Paolo, Durga, Chris R, Steve B, Paul K, Trevor H, Jaroslaw, Mike C, Victoria, John N, Paul H, Henry N, Ken L, Mark T, Colin W
Review of action items¶
- Realignment work
- Action Item (Colin to prod John C) : need to get John's data with fiducials in it to get to Tony Middleton. Expected yesterday
- Done
- Action Item (Colin to prod John C) : need to get John's data with fiducials in it to get to Tony Middleton. Expected yesterday
- Simulation of material in the interspace
- Action Item : John to do a simulation with air in the interspace and He in the absorber volume
- John presented plots to show that there is little effect. However, these plots include systematic effects in the reconstruction
that are being worked on. - Action Item : John will look at the effect of air on the MC truth scattering angle
- John presented plots to show that there is little effect. However, these plots include systematic effects in the reconstruction
- Action Item : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded
- Action Item : DC to work with Pierrick to ensure this
- Action Item : Colin will check with Alan Bross whether the LiH absorber can withstand vacuum
- Alan says that the LiH absorber can withstand vacuum.
- Action Item : John to do a simulation with air in the interspace and He in the absorber volume
- Cooling water discussion
- Action Item : Colin and John to talk and agree on the plan.
- Discussion complete but further work needed
- Action Item : Colin and John to talk and agree on the plan.
- ISIS Monitoring of SS magnets
- Action Item : Josef/Colin to check with Alan about monitoring requirements for the SS magnets.
- Josef doesn't think we need to monitor to the level we are doing now. Proposed to stop the 4-hour monitoring and just leave
the buzzer in the control room. - Action Item : Steve to talk to Alan Stevens about this possibility-
- Josef doesn't think we need to monitor to the level we are doing now. Proposed to stop the 4-hour monitoring and just leave
- Action Item : Josef/Colin to check with Alan about monitoring requirements for the SS magnets.
- BLOCS and SOCS
- Action Item : Steve to remind Adam that we need SOCs.
- Done
- Action Item : Steve to remind Adam that we need SOCs.
- DS PS Installation
- Action Item : Colin is to get an estimate for power supply installation time.
- 2 sets of 2-3 days over a space of 2 weeks.
- Action Item : Colin is to get an estimate for power supply installation time.
Updates and future schedule¶
- DS Power Supply work
- (Trevor) The PS is expected to return in mid-February (8th - 15th). He plans to start commissioning
the new PS on Feb 22nd. This could roll into the next week (which is nominally a data-taking week so may need to be postponed). - (Vicky) The compressor has a leak and will be fixed on Friday 29th and cleaned on Monday 1st
- (Discussion between Mike, Vicky and Henry) : The Linde IOC has been beahving oddly and we need to understand it.
- Action Item : Paolo, Mike and Vicky will investigate this issue ASAP. We can't cool down magnet until this is understood.
- (Discussion between Mike, Vicky and Henry) : We plan to start the cooldown of the cold box on Tuesday 2nd, with work on the controls and
monitoring doable whilst that is happening. If all is well, we can start the cooldown of the magnet on Thursday 4th. - (Discussion between Paolo and Mike) : we need a way of monitoring the status of the Linde IOC machine
- Action Item : Paolo/Pierrick to consider this issue
- (Trevor) The PS is expected to return in mid-February (8th - 15th). He plans to start commissioning
- Controls and Monitoring : Network reconfiguration
- (Paolo) Most work has been done but we wait for Pierrick to return to finish it. There is potentially another week of work with 1 - 2
days of severe disruption to the network whilst things are tested. This needs to be advertised well in advance
- (Paolo) Most work has been done but we wait for Pierrick to return to finish it. There is potentially another week of work with 1 - 2
- Realignment work
- Drilling of holes in the flanges will be done by Thursday 28th
- Template to show where the holes should go have been designed but need to be constructed.
- EMR will not be moved further at this time
- Revisit next week
- Tracker patch panel leak work
- John wants to take the front of the PRY off to get at the tracker leak, which he may have a solution for. This can be done now.
- DSA Work before closure on Friday 29th
- Need to reconnect the dipole magnets, and test all magnets in the DSA
- New drip trays will not be installed at this point
- Electrical work schedule
- (Trevor) Trevor/Steve Griffiths will be here on Monday 1st/Tuesday 2nd to do a low current test on the SS. This will
require the Hall closed up and the SS permit to be granted. DC will work with Trevor on this.- Action Item : (John G) need to put barriers around any open part of the SS
- Action Item : Andy/DC will organise the magnetic sweep
- A long discussion on the Daresbury scheduled ensued. We decided we needed a longer discussion.
- Action Item : Colin, Trevor, Steve, Pierrick to have a meeting to review the Daresbury work schedule in light of running time
and the currently known state of the SS work. Meeting to happen on morning of Thursday 28th.
- (Trevor) Trevor/Steve Griffiths will be here on Monday 1st/Tuesday 2nd to do a low current test on the SS. This will
Updated by Boyd, Steven over 7 years ago ยท 22 revisions